# Avoiding Catastrophe in Continuous Spaces by Asking for Help

### v2

@@ 2 Model
We will often deal with the measure of subsets of $S$. When $n=1$, we use $\operatorname*{len}(X)$ to denote the length of interval $X\subset S$; for general $n$, we use $\operatorname*{vol}(X)$ to denote the $n$-dimensional measure of $X\subseteq S$ [^7] Although we could use the same notation for both cases, we find this to be more intuitive. The convex hull of a set $X$ is denoted $\operatorname*{conv}(X)$. All logarithms and exponents are base $e$ unless otherwise noted.
3 Hardness result for complex mentor policies
We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when states are uniformly random on $[0,1]$ and there are only two actions. The result also holds even if the algorithm knows $L$ and $\mathbf{s}$ ahead of time.
Let $U(S)$ denote the uniform distribution on $S$ and let $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\mathbf{a}}$ denote the expectation over both $\mathbf{s}$ and any additional randomness in the algorithm.
@@ 3 Hardness result for complex mentor policies
The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity. Formally,
$\lim_{T\to\infty}\ \,\sup_{\mu,\pi^{*}}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ \,[R_{T}(\mathbf{s},\mu,\pi^{*})]=1$
@@ 3.1 Intuition
A natural idea is to try $f(T)=T$. Unfortunately, this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$. In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$. Lemma A.1 implies the existence of such an $f(T)$ if $|Q_{T}|$ is bounded by some function $g(T)\in o(T)$, so we get a bound of the following form: $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$. [^8] The bound in the actual proof will be in log space, so it will look slightly different, but the idea is the same. When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})=1$. In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
Figure 1: An illustration of the construction we use to prove Theorem 3.1 (not to scale). The horizontal axis indicates the state $s\in[0,1]$ and the vertical axis indicates the payoff $\mu(s,a)\in[0,1]$. The solid line represents $\mu(s,0)$ and the dotted line represents $\mu(s,1)$. In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by the Lipschitz constraint, reaching a minimum of $1-\frac{L}{2f(T)}$. Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section. As a result, we can freely toggle the optimal action for each section independently.
VC/Natarajan dimension.
The class of mentor policies induced by our construction has VC and Natarajan dimensions of $f(T)$. [^9] Recall that VC and Natarajan dimensions coincide for $|A|=2$. This is expected given our positive result for policy classes with bounded Natarajan dimension (Theorem 4.5).
@@ 3.2 Formal definition of construction
Let $S=[0,1]$, $A=\{0,1\}$, and $\mathcal{D}=U(S)$. Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent. We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\mathbf{x}=(x_{1},x_{2},\dots,x_{f(T)})\in\{0,1\}^{f(T)}$. The bit $x_{j}$ will denote the optimal action in section $j$. Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
@@
For any $\mathbf{x}\in\{0,1\}^{f(T)}$, $\mu_{f,\mathbf{x}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections). Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\mathbf{x}}$ is Lipschitz continuous with Lipschitz constant $L$.
@@
Since each $s_{t}$ belongs to exactly one $S_{j}$, there is a unique optimal action for each $s_{t}$, so the choice of $\pi^{*}$ does not matter in this scenario.
4 Near-zero regret assuming bounded Natarajan dimension
In this section, we present a simple algorithm with near-zero regret when the policy class $\Pi$ has bounded Natarajan dimension $d$ and which contains at least some “reasonable” policies. The algorithm needs to know $\Pi$ and (an upper bound on) $d$, as is standard in PAC learning (see, e.g., [27]). The algorithm does not need to know $L$, and can handle an unbounded state space (the number of queries simply scales with the volume of the convex hull of the observed states).
We need a few definitions to formally state our algorithm and result.
@@ 4 Near-zero regret assuming bounded Natarajan dimension
A policy $\pi$ is consistent with $X\subset S\times A$ if $\pi(s)=a$ for all $(s,a)\in X$.
@@
Given $X\subset S\times A$ and $r\geq 0$, a policy $\pi$ is $(X,r)$-local if (1) $\pi$ is consistent with $X$ and (2) whenever $(s^{\prime},a^{\prime})\in X$ and $||s-s^{\prime}||\leq r$, there exists $(s^{\prime\prime},a^{\prime\prime})\in X$ such that $||s-s^{\prime\prime}||\leq r$ and $\pi(s)=a^{\prime\prime}$.
A policy which knows $X$ but is not $(X,r)$-local says either (1) “even though I know the optimal action for $s$, I’m going to do something else” or (2) “even though I have a data point near $s$, I’m going to base my action on a data point far from $s$”. [^10] Note that the mentor policy cannot itself query the mentor, so we are only dealing with actions in $A$. Given the Lipschitz nature of the problem, this type of policy seems quite unintuitive. However, we allow $\Pi$ to contain non-local policies, as long as it also contains local policies. Note that knowing $X,r$, and $\Pi$ is sufficient to find a $(X,r)$-local policy.
Definition 4.3.
Given $r\geq 0$, $\Pi$ is $r$-local if for every $\tilde{S}\subset S$ and $X=\{(s,\pi^{*}(s)):s\in\tilde{S}\}$, there exists an $(X,r)$-local policy in $\Pi$.
As an example, let $S=\mathbb{R}$ and suppose $\Pi$ is the set of policies where $\pi(s)$ changes at most $K$ times as $s$ goes from $\infty$- ∞ to +$\infty$+ ∞ (in the language of Section 5, $\pi$ has at most $K$ segments). Then $\Pi$ is $r$-local for any $r$, since for any set of queries $X$, the nearest neighbor policy is $(X,r)$-local and has at most $K$ segments (and thus is in $\Pi$). Note that this $\Pi$ also has bounded Natarajan dimension.
4.1 Intuition behind the algorithm
Our algorithm has a simple structure: gather enough data to choose a solid baseline policy, and then mostly follow that policy, but ask for help if you find yourself out-of-distribution (OOD). The algorithm does not need direct access to the feature embedding of states: the only two operations it uses are (1) update baseline policy and (2) find nearest neighbor distance. This approach seems quite natural for real-world applications, where one can imagine updating the baseline policy via gradient descent and treating an OOD detector score as the nearest neighbor distance.
In our case, the initial dataset $X$ is gathered by querying on the first $g(T)$ time steps. We will choose $g(T)$ to be large enough so that – along with the assumption on Natarajan dimension – we can apply the fundamental theorem of multi-class PAC learning (Theorem 4.4). [^11] Theorem 4.4 is called the “realizable” case of PAC learning, in contrast to the “agnostic” case where $\pi^{*}$ may not be in $\Pi$. Some sources (e.g., [27]) state that $\pi$ should minimize empirical loss, rather than being consistent with $X$. These are equivalent in the realizable case since consistent policies have zero empirical loss. This is the only place in the proof where Natarajan dimension is used.
Theorem 4.4 (Fundamental Theorem of Multi-Class PAC Learning [13; 24; 27]).
Assume $\Pi$ has finite Natarajan dimension $d$ and $\pi^{*}\in\Pi$. Then $\exists\,c>0$ such that $\forall\,\varepsilon,\delta\in(0,1)$, if $X\subset S\times A$ consists of at least $\frac{c}{\varepsilon}\left(d\log(\frac{|A|d}{\varepsilon})+\log\frac{1}{\delta}\right)$ i.i.d. samples $\mathcal{D}$ and the corresponding actions from $\pi^{*}$, then with probability at least $1-\delta$, any policy $\pi\in\Pi$ consistent with $X$ satisfies $\Pr_{s\sim\mathcal{D}}[\pi(s)\neq\pi^{*}(s)]\leq\varepsilon$.
Theorem 4.4 will ensure that $\Pr[a_{t}\neq\pi^{*}(s_{t})]$ is small, but even a single incorrect action can spell defeat if the payoff is bad enough. To handle this, the algorithm continues to selectively ask for help, but only when the state is far from any previously queried state (i.e., out-of-distribution). The policy $\pi$ is then updated based on this new information. The threshold $r$ is chosen such that we can cover the entire state $S$ with a sublinear number of queries.
1:function LookBeforeYouLeap($T\in\mathbb{N},\>r\in\mathbb{R}_{\geq 0},\>g:\mathbb{N}\to\mathbb{N}$)
2:     $X\leftarrow\emptyset$
3:     for $t$ from $1$ to $g(T)$ do $\triangleright$ Gather baseline data set
4:         Query mentor and observe $\pi^{*}(s_{t})$
5:         $X\leftarrow X\cup\{(s_{t},\pi^{*}(s_{t}))\}$
6:     $\pi\leftarrow$ any $(X,r)$-local policy in $\Pi$
7:     for $t$ from $g(T)+1$ to $T$ do $\triangleright$ Mostly follow $\pi$, but ask for help if out-of-distribution
8:         if $\min_{(s,a)\in X}||s_{t}-s||\leq r$ then
9:              Take action $\pi(s_{t})$
10:         else
11:              Query mentor and observe $\pi^{*}(s_{t})$
12:              $X\leftarrow X\cup\{(s_{t},\pi^{*}(s_{t}))\}$
13:              $\pi\leftarrow$ any $(X,r)$-local policy in $\Pi$
Algorithm 1 achieves near-zero regret when $\Pi$ has bounded Natarajan dimension and is $r$-local.
Tractability. Our algorithm can only be useful in practice if finding a $(X,r)$-local policy in $\Pi$ is computationally tractable. Such guarantees are impossible without further assumptions on $\Pi$ (this is again standard for PAC learning). However, if $\Pi$ supports nearest-neighbor policies (which satisfy locality for all values for $r$), those policies are both intuitive and can be computed easily.
We are now ready to state the main result of this section. The constant $c$ in Theorem 4.5 is the same as the constant $c$ in Theorem 4.4.
Theorem 4.5.
[…]

### v3

@@ 3 Model
The diameter of $X\subseteq S$ is defined by $\operatorname*{diam}(X)=\max_{x,y\in X}||x-y||$. All logarithms and exponents are base $e$ unless otherwise noted.
4 Avoiding catastrophe is impossible in general
We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when states are i.i.d. on $[0,1]$. The result also holds even if the algorithm knows $L$ and $\mathbf{s}$ ahead of time. We use $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}$ denote the expectation over both $\mathbf{s}$ and any additional randomness in the algorithm.
@@ 4 Avoiding catastrophe is impossible in general
The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity. Formally, $\lim_{T\to\infty}\ \,\sup_{\mu,\pi^{m}}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}\ \,[R_{T}]=1$.
@@ 4.1 Intuition
Figure 3: An illustration of the construction we use to prove Theorem 4.1 (not to scale). The horizontal axis indicates the state $s\in[0,1]$ and the vertical axis indicates the payoff $\mu(s,a)\in[0,1]$. The solid line represents $\mu(s,0)$ and the dotted line represents $\mu(s,1)$. In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by $L$, reaching a minimum of $1-\frac{L}{2f(T)}$. Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section. As a result, we can freely toggle the optimal action for each section independently.
@@
A natural idea is to try $f(T)=T$, but this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$. In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$ (such $f$ must exist since $|Q_{T}|\leq g(T)\in o(T)$). This leads to the following bound: $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$. When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu^{m}(s_{t})=1$. In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
VC dimension.
The class of mentor policies induced by our construction has VC dimension $f(T)$; considered over all possible values of $T$, this implies infinite VC (and Littlestone) dimension. This is necessary given our positive results in Section 5.
@@ 4.2 Formal definition of construction
Let $S=[0,1]$ and $\mathcal{D}_{t}=U$ for each $t\in[T]$. Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent. We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\mathbf{x}=(x_{1},x_{2},\dots,x_{f(T)})\in\{0,1\}^{f(T)}$. The bit $x_{j}$ will denote the optimal action in section $j$. Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
@@
Let $\pi^{m}$ be any policy which is optimal for $\mu_{f,\mathbf{x}}$. Note that there is a unique optimal action for each $s_{t}$, since each $s_{t}$ belongs to exactly one $S_{j}$; formally, $\pi^{m}(s_{t})=x_{j(s_{t})}$.
For any $\mathbf{x}\in\{0,1\}^{f(T)}$, $\mu_{f,\mathbf{x}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections). Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\mathbf{x}}$ is Lipschitz continuous. Thus by Proposition E.1, $\pi^{m}$ satisfies local generalization.
5 Avoiding catastrophe given finite VC or Littlestone dimension
Theorem 4.1 shows that avoiding catastrophe is impossible in general, which is also true in online learning. What if we restrict ourselves to settings where standard online learning is possible? Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^8] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
This section presents a simple algorithm which guarantees subconstant regret and sublinear queries under either of those assumptions. Our algorithm needs to know $\Pi$, as is standard in online learning. The algorithm does not need to know $\sigma$ (in the smooth case) or $L$, and can handle an unbounded state space (the number of queries simply scales with the maximum distance between observed states). For simplicity, we initially prove our result for $A=\{0,1\}$. Appendix C extends our result to many actions using the“one versus rest” reduction. [^9] For each action $a$, we learn a binary classifier which predicts whether $\pi^{m}(s)=a$. If every binary classifier is correct, we can correctly determine $\pi^{m}(s)$. See, e.g., Chapter 29 of [34].
5.1 Intuition behind the algorithm
Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help in unfamiliar states (specifically, when the current state is far from any queried state with the same action under the proposed policy). Hedge ensures that the number of time steps where the agent’s action doesn’t match the mentor’s is small, and asking for help in unfamiliar states ensures that when we do make a mistake, the cost isn’t too high. This algorithmic structure seems quite natural: mostly follow a baseline strategy, but ask for help when out-of-distribution.
Simple operations.
The algorithm does not require detailed access to the state embedding, instead relying on two simple operations: evaluating a policy on a particular state, and computing a nearest neighbor distance. The former seems necessary for any algorithm. The latter could be modeled as an out-of-distribution detector score, for which many methods are available (see e.g., [41]).
Hedge.
Hedge [12] is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite. [^10] See Chapter 5 of [35] and Chapter 21 of [34] for modern introductions to Hedge. We would prefer not to assume that $\Pi$ is finite. Luckily, any policy $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\mathbf{s}$ $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension. Thus we can run Hedge on this approximative policy class instead.
One other modification is necessary. In standard online learning, losses are observed on every time step, but our agent only receives feedback in response to queries. To handle this, we modify Hedge to only perform updates on time steps with queries and to issue a query with probability $p$ on each time step. Continuing our lucky streak, [33] analyzes exactly this modification of Hedge.
1:function AvoidCatastrophe($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
2:     if $\Pi$ has VC dimension $d$ then
3:         $\tilde{\Pi}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$triangleright See Definition 5.3
4:     else
5:         $\tilde{\Pi}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$triangleright See Definition 5.4
6:     $X\leftarrow\emptyset$
7:     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
8:     $p\leftarrow 1/\sqrt{\varepsilon T}$
9:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
10:     for $t$ from $1$ to $T$ do $\triangleright$ Run one step of Hedge, which selects policy $\pi_{t}$
[…]
12:         if hedgeQuery then
13:              Query mentor and observe $\pi^{m}(s_{t})$
14:              $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$ for all $\pi\in\tilde{\Pi}$
15:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
16:              $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
17:              $\pi_{t}\leftarrow\operatorname*{arg\,min}_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
18:         else
19:              $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
20:              Sample $\pi_{t}\sim P$
21:         if $\min_{(s,a)\in X:a=\pi_{t}(s_{t})}||s_{t}-s||>\varepsilon^{1/n}$ then $\triangleright$ Ask for help if out-of-distribution
22:              Query mentor and observe $\pi^{m}(s_{t})$
23:              $X\leftarrow X\cup\{(s_{t},\pi^{m}(s_{t}))\}$
24:         $\triangleright$ Otherwise, follow Hedge’s chosen policy
25:              Take action $\pi_{t}(s_{t})$
Algorithm 1 successfully avoids catastrophe assuming finite VC or Littlestone dimension.
We prove the following theorem parametrized by $\varepsilon$:
Theorem 5.1.
Let $A=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$, $\mathbf{s}$ is $\sigma$-smooth, and $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$ and $\varepsilon>0$, Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
In Case 1, the expectation is over the randomness of both $\mathbf{s}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm. Also, $R_{T}$ and $Q_{T}$ clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate bounds.
To obtain subconstant regret and sublinear queries, we can choose $\varepsilon=T^{\frac{-2n}{2n+1}}$. This also satisfies the requirement of $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ for large enough $T$.
Theorem 5.2.
Let $A=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)$
We can also make the query bound unconditional in exchange for slightly worse bounds. The probability that $|Q_{T}|$ asymptotically exceeds the bound from Theorem 5.2 goes to 0, so the query threshold in Corollary 5.2.1 is almost never reached. Thus we still obtain subconstant regret.
Corollary 5.2.1.
If Algorithm 1 is modified to stop querying after $T^{\frac{4n+1.5}{4n+2}}(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n})$ queries, then under the conditions of Theorem 5.2, we have $\operatorname*{\mathbb{E}}[R_{T}]\in O\big{(}\frac{dL}{\sigma}T^{\frac{-1}{8n+4}}\log T\big{)}$.
5.2 Proof sketch
The formal proof of Theorem 5.1 can be found in Appendix B, but we outline the key elements here. The regret analysis consists of two ingredients: analyzing the Hedge component, and analyzing the “ask for help when out-of-distrubtion” component. The former will bound the number of mistakes made by the algorithm (i.e., the number of time steps where the agent’s action doesn’t match the mentor’s), and the latter will bound the cost of any single mistake. We must also carefully show that the latter does not result in excessively many queries, which we do via a novel packing argument.
We begin by formalizing two notion of approximating a policy class:
@@ 5.2 Proof sketch
Let $U$ be the uniform distribution over $S$. For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ is for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{s\sim U}[\pi(s)\neq\tilde{\pi}(s)]\leq\varepsilon$.
@@
A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ is for every $\mathbf{s}\in S^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(s_{t})=\tilde{\pi}(s_{t})$ for all $t\in[T]$.
The existence of small covers is crucial:
Lemma 5.1 (Lemma 7.3.2 in [16] [^11] See also [19] or Lemma 13.6 in [6] for variants which are less convenient for our purposes.).
For all $\varepsilon>0$, any policy class of VC dimension $d$ admits a smooth $\varepsilon$-cover of size at most $(41/\varepsilon)^{d}$.
Lemma 5.2 (Lemmas 21.13 and A.5 in [34]).
Any policy class of Littlestone dimension $d$ admits an adversarial cover of size at most $(eT/d)^{d}$.
An adversarial cover is a perfect cover by definition. The following lemma establishes that a smooth $\varepsilon$-cover is a good approximation for any sequence of $\sigma$-smooth distributions.
Lemma 5.3 (Equation 2 and Lemma 3.3 in [17]).
Let $\tilde{\Pi}$ be a finite smooth $\varepsilon$-cover of $\Pi$ and let $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ be a sequence of $\sigma$-smooth distributions. If $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$, then